OpenStudy (anonymous):

(b -c) ______ (2b - a)

OpenStudy (loser66):

What are we supposed to do?

OpenStudy (anonymous):

i don't know where to started

OpenStudy (anonymous):

This expression uses all three variables: a, b, and c. It is a rational expression with a divisor of 2b – a. The integers have been plugged in for the variables. Both numerator and denominator have been evaluated. Both are negative which means the answer will be positive. This is our answer in lowest terms, and I will leave the answer as an improper fraction.

OpenStudy (loser66):

I'm sorry for being dummy. I don't get what you mean, still. :( Ok, let me try to interpret it if a =1, b=2, c =3 then the expression is \(\dfrac{b-c}{2b-a}=\dfrac{2-3}{2*2-1}=\dfrac{-1}{3}\)

OpenStudy (anonymous):

i have the diection

OpenStudy (anonymous):

yes that it

OpenStudy (anonymous):

but i have to use my birthday to replace

OpenStudy (anonymous):

a= 4 b=-30 c= 85